# Pay Attention to Your Neighbours: Training-Free Open-Vocabulary Semantic Segmentation

### v1

@@ 6 Conclusion
Drawing from CLIP’s remarkable zero-shot generalizability, the paradigm of open-vocabulary semantic segmentation leveraging this model has gained significant traction, emerging as an appealing alternative to circumvent the limitations of traditional close-set supervised training. In this work, we have explored the inherent weaknesses of CLIP for dense prediction and proposed simple and minimal modifications that accommodate this powerful model to the restrictive training-free OVSS scenario. In addition to removing components of CLIP’s visual encoder that hamper its localization capabilities, we have integrated a simple mechanism that explicitly encourages local consistency in the self-attention maps, which has been unexplored in the existing works. Extensive experiments on popular OVSS benchmarks showcase the superiority of our approach over other existing OVSS methods, some of which resort to impractical or unrealistic choices, such as leveraging auxiliary models trained on additional large datasets or relying on validation sets for hyperparameter tuning. In contrast, NACLIP does not require access to either labeled or unlabeled data, becoming a suitable solution for real-world applications.
6.0.1 Limitations.
In CLIP’s pre-training, only the output at the [CLS] token’s position directly influences optimization [27]. However, given the fundamental disparity between CLIP’s pre-training objective and the task of semantic segmentation, we deemed the [CLS] token potentially irrelevant for dense prediction and consequently excluded it from our process. Acknowledging the proven efficacy of the [CLS] token’s output representation in various image-level tasks, we recognize the possibility of its applicability to dense prediction tasks as well. Hence, our inability to leverage the [CLS] token represents a limitation necessitating further investigation and assessment.
@@ Acknowledgments
Appendix 0.A Additional Comparison with SCLIP
In Tab. 1, we included the performance of NACLIP with and without PAMR [1]. Additionally, we clarified that for methods that opt to use post-processing steps, such as SCLIP [32], reported results are based on their enhanced performance after the post-processing step. Given SCLIP’s significance as a primary competitor, we conduct a comparative analysis in this section, examining the performance of both methods with and without PAMR (refer to Tab. 5) to isolate the actual performance of each approach. These results demonstrate our method’s superiority over SCLIP in both scenarios.
Table 5: Comparison of NACLIP with SCLIP [32] across various benchmarks. We present results for both methods with and without the post-processing mechanism, i.e., PAMR [1]. The abbreviations for dataset names are explained in Sec. 5.1.
Method PAMR V21 PC60 C-Obj V20 City PC59 ADE C-Stf Avg SCLIP [32] 59.1 30.4 30.5 80.4 32.2 34.2 16.1 22.4 38.2 NACLIP 58.8 32.2 33.2 77.1 35.5 35.2 17.4 22.9 39.0 SCLIP [32] 61.7 31.5 32.1 83.5 34.1 36.1 17.8 23.9 40.1 NACLIP 62.4 35.0 36.2 80.6 38.3 38.4 19.1 25.2 41.9
Appendix 0.B Gaussian Kernel’s Standard Deviation
@@
For a patch located at $\boldsymbol{\mu}$, the Gaussian kernel increases its attention logits by $1$ at $\boldsymbol{\mu}$ and by lesser values at neighbouring patch locations. Our choice of $\sigma$ is based on the count of neighbouring patches whose attention logits are modified by more than a threshold $\tau$. To achieve this, we express:
@@
Table 6: Proposed heuristic measure for determining $\sigma$ value. For 3 values of $\tau$ and 10 values of $\sigma$, the table provides the number of patches that patch $\boldsymbol{\mu}$’s attention logits to them is increased by more than $\tau$. It is important to note that these values are derived based on an infinite grid of patches, while in practice, these numbers could be less depending on the window size and the position of $\boldsymbol{\mu}$.
@@
Although we employed a heuristic measure to determine $\sigma$, we provide in Fig. 4 the impact of varying $\sigma$ values on test set performance. Please note that these experiments were conducted after deciding to use $\sigma=5$, and whose goal is to demonstrate: i) that our heuristic approach to find $\sigma$ provides indeed a good value; and ii) that the performance across different datasets is not strongly sensitive to the hyperparameter $\sigma$.
@@
Appendix 0.C Visual Examples
@@
Figure 5: Visual examples from PASCAL Context (59) [25]. Segmentation maps of vanilla CLIP [27], SCLIP [32], and our method are presented.
Figure 6: Additional visual examples from COCO-Object [20, 5]. Segmentation maps of vanilla CLIP [27], SCLIP [32], and our method are presented.

### v2

@@ 6 Conclusion
Drawing from CLIP’s remarkable zero-shot generalizability, the paradigm of open-vocabulary semantic segmentation leveraging this model has gained significant traction, emerging as an appealing alternative to circumvent the limitations of traditional close-set supervised training. In this work, we have explored the inherent weaknesses of CLIP for dense prediction and proposed simple and minimal modifications that accommodate this powerful model to the restrictive training-free OVSS scenario. In addition to removing components of CLIP’s visual encoder that hamper its localization capabilities, we have integrated a simple mechanism that explicitly encourages local consistency in the self-attention maps, which has been unexplored in the existing works. Extensive experiments on popular OVSS benchmarks showcase the superiority of our approach over other existing OVSS methods, some of which resort to impractical or unrealistic choices, such as leveraging auxiliary models trained on additional large datasets or relying on validation sets for hyperparameter tuning. While yielding state-of-the-art performance on 7 out of 8 benchmarks, NACLIP does not require access to either labeled or unlabeled data, becoming a suitable solution for real-world applications.
Limitations.
In CLIP’s pre-training, only the output at the [CLS] token’s position directly influences optimization [32]. Although this token plays a primary role in the original CLIP’s encoders, our attempts to extract information useful for segmentation from it were unsuccessful, having almost no effect on performance. We argue that this failure can be attributed to the divergent nature of CLIP’s pre-training and the specific requirements of dense prediction problems, in which objects of interest must be recognized with their local positioning. Despite this, we acknowledge that the [CLS] token has proven effective in many image-level tasks and may contain information transferable to dense prediction. Therefore, we believe the use of this token in dense prediction requires further investigation.
@@ Acknowledgments
Supplementary Material for
Pay Attention to Your Neighbours:
Training-Free Open-Vocabulary Semantic Segmentation
Appendix A Gaussian kernel’s standard deviation
@@
For a patch located at $\boldsymbol{\mu}$, the Gaussian kernel increases its attention logits by $1$ at $\boldsymbol{\mu}$ and by lesser values at neighbouring patch locations. Our choice of $\sigma$ is based on the number of neighbouring patches whose attention logits are modified by more than a threshold $\tau$. To achieve this, we express:
@@
Table 5: Proposed heuristic measure for determining $\sigma$ value. For 3 values of $\tau$ and 10 values of $\sigma$, the table provides the number of patches that patch $\boldsymbol{\mu}$’s attention logits to them is increased by more than $\tau$. It is important to note that these values are derived based on an infinite grid of patches, while in practice, these numbers could be less depending on the window size and $\boldsymbol{\mu}$’s position.
@@
Although we employed a heuristic measure to determine $\sigma$, we provide in Fig. 5 the impact of varying $\sigma$ values on test set performance. Please note that these experiments were conducted after deciding to use $\sigma=5$, and whose goal is to demonstrate that i) our heuristic approach to find $\sigma$ provides indeed a good value; and ii) the performance across different datasets is not strongly sensitive to the hyperparameter $\sigma$.
@@
Appendix B Visual examples
@@
Figure 6: Additional visual examples (segmentation maps) from PASCAL Context (59) [29] for CLIP [32], SCLIP [38], and our method.
Figure 7: Additional visual examples (segmentation maps) from COCO-Object [23, 5] for CLIP [32], SCLIP [38], and our method.